OpenStudy (anonymous):

You have two block of clay in cube form and the edges are 10 cm. How many spheres with a radius of 5 cm can you make with that amount of clay?

OpenStudy (anonymous):

Volume of sphere = (4/3)pi*r^3

OpenStudy (anonymous):

Volume of cube = a^3 where a is the side of the cube

OpenStudy (anonymous):

So. The answer is 3?

OpenStudy (callisto):

volume of cubes available = 10^3 x2= 2000 cm^3 volume of each sphere = 4/3 (pi) (5)^3 = 523.59878cm^3 no. of sphere can be made = 2000/523.59878 = 3.81, that is 3 spheres can be made

OpenStudy (anonymous):

So, 3 is right. @Callisto You know, you could of just said that, but thanks anyway.

OpenStudy (callisto):

lol just saying the answer will be reported, i received a warning few hours ago .. it's a place to teach how to get the answer, not to give the answer

OpenStudy (anonymous):

-.-' That's just sad.

OpenStudy (anonymous):

Because I just wanted to check if my answer were right -.-'

OpenStudy (callisto):

perhaps you can also check the workings :)

OpenStudy (anonymous):

Why would I? It's just waste of time.